プリンキピア

付録

4.5: Hypergeometric Distribution4.5:超幾何分布

35. Yes, because you are sampling from a population composed of two groups (boys and girls), have a group of interest (boys), and are sampling without replacement (hence, the probabilities change with each pick, and you are not performing Bernoulli trials).
35。はい。二つの集団(男子と女子)からなる母集団から抽出し、注目する集団(男子)があり、非復元抽出を行うためである。したがって、引くたびに確率は変わり、ベルヌーイ試行ではない。
36. The group of interest is the cards that are spades, the size of the group of interest is 13, and the sample size is five.
36。注目する集団はスペードのカードで、その大きさは13、標本の大きさは5である。

4.6: Poisson Distribution4.6:ポアソン分布

37. A Poisson distribution models the number of events occurring in a fixed interval of time or space, when the events are independent and the average rate of the events is known.
37。ポアソン分布は、事象が独立で、平均発生率がわかっているとき、一定の時間または空間の区間で起きる事象の数をモデル化する。
38. X ~ P(4)
38。X ~ P(4)
39. The domain of X = {0, 1, 2, 3, …..) i.e., any integer from 0 upwards.
39。X の取りうる値は {0, 1, 2, 3, …}、すなわち0以上の任意の整数である。
40. μ = 4
σ = 4 = 2

5.1: Continuous Probability Functions5.1:連続型確率関数

41. The discrete variables are the number of books purchased, and the number of books sold after the end of the semester. The continuous variables are the amount of money spent for the books, and the amount of money received when they were sold.
41。離散型の変数は購入した冊数と学期終了後に売った冊数である。連続型の変数は購入額と売却時に受け取った金額である。
42. Because for a continuous random variable, P(x = c) = 0, where c is any single value. Instead, we calculate P(c < x < d), i.e., the probability that the value of x is between the values c and d.
42。連続型確率変数では、任意の特定の値 c について P(x = c) = 0 となるためである。代わりに P(c < x < d)、すなわち x が c と d の間にある確率を計算する。
43. Because P(x = c) = 0 for any continuous random variable.
43。どの連続型確率変数でも P(x = c) = 0 となるため。
44. P(x > 5) = 1 – 0.35 = 0.65, because the total probability of a continuous probability function is always 1.
44。連続型確率関数の確率の合計は常に1なので、P(x > 5) = 1 − 0.35 = 0.65 である。
45. This is a uniform probability distribution. You would draw it as a rectangle with the vertical sides at 0 and 20, and the horizontal sides at 1 10 and 0.
45。これは一様確率分布である。縦の辺を0と20、横の辺を 1 10 と0に置く長方形として描く。
46. P ( 0 < x < 4 ) = ( 4 − 0 ) ( 1 10 ) = 0.4

5.2: The Uniform Distribution5.2:一様分布

47. P ( 2 < x < 5 ) = ( 5 − 2 ) ( 1 10 ) = 0.3
48. X ~ U(0, 15)
48。X ~ U(0, 15)
49. f ( x ) = 1 b − a for ( a ≤ x ≤ b ) so f ( x ) = 1 30 for ( 0 ≤ x ≤ 30 )
49。 f ( x ) = 1 b − a 、範囲は ( a ≤ x ≤ b ) したがって f ( x ) = 1 30 、 ( 0 ≤ x ≤ 30 ) である。
50. μ = a + b 2 = 0 + 30 5 = 15.0
σ = ( b − a ) 2 12 = ( 30 − 0 ) 2 12 = 8.66
51. P ( x < 10 ) = ( 10 ) ( 1 30 ) = 0.33

5.3: The Exponential Distribution5.3:指数分布

52. X has an exponential distribution with decay parameter m and mean and standard deviation 1 m . In this distribution, there will be a relatively large numbers of small values, with values becoming less common as they become larger.
英語のヒント
52。X は、減衰パラメータ m、平均と標準偏差 1 m の指数分布に従う。この分布には小さい値が比較的多く、値が大きくなるほど少なくなる。
53. μ = σ = 1 m = 1 10 = 0.1
54. f(x) = 0.2e–0.2x where x ≥ 0.
54。x ≥ 0 に対して、f(x) = 0.2e–0.2x。

6.1: The Standard Normal Distribution6.1:標準正規分布

55. The random variable X has a normal distribution with a mean of 100 and a standard deviation of 15.
55。確率変数 X は、平均100、標準偏差15の正規分布に従う。
56. X ~ N(0,1)
56。X ~ N(0, 1)
57. z = x − μ σ so z = 112 − 109 4.5 = 0.67
57。 z = x − μ σ より、 z = 112 − 109 4.5 = 0.67 。
58. z = x − μ σ so z = 100 − 109 4.5 = − 2.00
58。 z = x − μ σ より、 z = 100 − 109 4.5 = − 2.00 。
59. z = 105 − 109 4.5 = −0.89
This girl is shorter than average for her age, by 0.89 standard deviations.
59。 z = 105 − 109 4.5 = −0.89 この女子の身長は、同年齢の平均より標準偏差の0.89倍だけ低い。
60. 109 + (1.5)(4.5) = 115.75 cm
60。109 + (1.5)(4.5) = 115.75 cm
61. We expect about 68 percent of the heights of girls of age five years and zero months to be between 104.5 cm and 113.5 cm.
61。5歳0か月の女子の身長の約68%が、104.5~113.5cmに入ると予想される。
62. We expect 99.7 percent of the heights in this distribution to be between 95.5 cm and 122.5 cm, because that range represents the values three standard deviations above and below the mean.
62。95.5~122.5cmは平均から上下に標準偏差の3倍の範囲なので、この分布の身長の99.7%が入ると予想される。

6.2: Using the Normal Distribution6.2:正規分布の利用

63. Yes, because both np and nq are greater than five.
np = (500)(0.20) = 100 and nq = 500(0.80) = 400
63。はい。np と nq がどちらも5より大きいため。np = (500)(0.20) = 100、nq = 500(0.80) = 400。
64. μ = n p = ( 500 ) ( 0.20 ) = 100
σ = n p q = 500 ( 0.20 ) ( 0.80 ) = 8.94
65. Fifty percent, because in a normal distribution, half the values lie above the mean.
65。50%。正規分布では、半数の値が平均より上にあるため。
66. The results of our sample were two standard deviations below the mean, suggesting it is unlikely that 20 percent of the lotto tickets are winners, as claimed by the distributor, and that the true percent of winners is lower. Applying the Empirical Rule, If that claim were true, we would expect to see a result this far below the mean only about 2.5 percent of the time.
66。標本の結果は平均より標準偏差の2倍だけ低かった。これは、販売者が主張するように券の20%が当たりである可能性は低く、実際の当たりの割合はもっと低いことを示唆する。経験則によると、主張が正しければ、平均をこれほど下回る結果が出るのは約2.5%の場合だけと予想される。

7.1: The Central Limit Theorem for Sample Means (Averages)7.1:標本平均についての中心極限定理

67. The central limit theorem states that if samples of sufficient size drawn from a population, the distribution of sample means will be normal, even if the distribution of the population is not normal.
67。中心極限定理は、母集団から十分な大きさの標本を取れば、母集団が正規分布でなくても、標本平均の分布は正規分布になると述べる。
68. The sample size of 30 is sufficiently large in this example to apply the central limit theorem. This theorem ] states that for samples of sufficient size drawn from a population, the sampling distribution of the sample mean will approach normality, regardless of the distribution of the population from which the samples were drawn.
68。この例では標本の大きさ30は、中心極限定理を適用できるほど大きい。定理は、母集団から十分な大きさの標本を取ると、元の母集団の分布にかかわらず、標本平均の標本分布は正規分布に近づくと述べる。
69. You would not expect each sample to have a mean of 50, because of sampling variability. However, you would expect the sampling distribution of the sample means to cluster around 50, with an approximately normal distribution, so that values close to 50 are more common than values further removed from 50.
69。標本抽出の変動があるため、すべての標本の平均が50になるとは期待できない。しかし、標本平均の標本分布は50の周りに集まり、およそ正規分布になると期待できる。そのため50に近い値は多く、50から離れた値は少なくなる。
70. X ¯ ∼ N ( 25 , 0.2 ) because X ¯ ∼ N ( μ x , σ x n )
70。 X ¯ ∼ N ( 25 , 0.2 ) 。 X ¯ ∼ N ( μ x , σ x n ) だからである。
71. The standard deviation of the sampling distribution of the sample means can be calculated using the formula ( σ x n ) , which in this case is ( 16 50 ) . The correct value for the standard deviation of the sampling distribution of the sample means is therefore 2.26.
71。標本平均の標本分布の標準偏差は、公式 ( σ x n ) で計算でき、この場合は ( 16 50 ) である。したがって、正しい標準偏差は2.26である。
72. The standard error of the mean is another name for the standard deviation of the sampling distribution of the sample mean. Given samples of size n drawn from a population with standard deviation σx, the standard error of the mean is ( σ x n ) .
72。平均の標準誤差は、標本平均の標本分布の標準偏差の別名である。標準偏差 σx の母集団から大きさ n の標本を取るとき、平均の標準誤差は ( σ x n ) である。
73. X ~ N(75, 0.45)
73。X ~ N(75, 0.45)
74. Your friend forgot to divide the standard deviation by the square root of n.
74。友人は、標準偏差を n の平方根で割り忘れている。
75. z = x ¯ − μ x σ x = 76 − 75 4.5 = 2.2
76. z = x ¯ − μ x σ x = 74.7 − 75 4.5 = −0.67
77. 75 + (1.5)(0.45) = 75.675
78. The standard error of the mean will be larger, because you will be dividing by a smaller number. The standard error of the mean for samples of size n = 50 is:
( σ x n ) = 4.5 50 = 0.64
78。割る数が小さくなるため、平均の標準誤差は大きくなる。n = 50 の標本の平均の標準誤差は、 ( σ x n ) = 4.5 50 = 0.64 。
79. You would expect this range to include values up to one standard deviation above or below the mean of the sample means. In this case:
70 + 9 60 = 71.16 and 70 − 9 60 = 68.84 so you would expect 68 percent of the sample means to be between 68.84 and 71.16.
79。標本平均の平均から上下に標準偏差一つ分までの値が、この範囲に入ると予想される。この場合、 70 + 9 60 = 71.16 、 70 − 9 60 = 68.84 なので、標本平均の68%は68.84~71.16の間に入ると予想される。
80. 70 + 9 100 = 70.9 and 70 − 9 100 = 69.1 so you would expect 68 percent of the sample means to be between 69.1 and 70.9. Note that this is a narrower interval due to the increased sample size.
80。 70 + 9 100 = 70.9 、 70 − 9 100 = 69.1 なので、標本平均の68%は69.1~70.9の間に入ると予想される。標本の大きさが増えたため、区間は狭くなっている。